# Emergent predictability in microbial ecosystems

### v2

@@ III Discussion
In this work, we developed an information-theoretic framework to nuance and quantify the notion of “emergent simplicity” in ecosystems of increasing community richness (treated as the control parameter). Our framework enabled us to define and distinguish emergent predictability and emergent coarse-grainability, and to contrast both of these phenomena with the emergent reproducibility considered in much previous work. We applied our framework to data from two separate experiments on synthetic microbial communities, and demonstrated that four out of five measured properties exhibit robust evidence of emergent predictability, whereby simple compositional descriptions become more predictive at higher richness.
Figure 5: Shifting the focus from simple and complex ecosystems to simple and complex properties of ecosystems. As demonstrated in this work, even in a complex (high-richness) community, some properties may be adequately predicted with a low-complexity description; we call such properties coarse-grainable (green Pareto front).
In three of these four cases, the phenomenon appears to be linked to the dominant structuring role of pH in these communities. This observation echoes other recent examples where coarse-grained predictions are enabled by general physiological [10], metabolic [3], or toxicity-mediated constraints [35]. Our results indicate that these structuring factors may have higher explanatory power at higher diversity, suggesting the existence of a subtler form of functional convergence, where high diversity leads not to better absolute reproducibility, but to a stronger collapse onto a lower-dimensional subspace governed by metabolic or physiological constraints. Understanding this behavior will require integrating ongoing theoretical investigations of large random ecosystems with new models capable of incorporating the structuring role of such shared constraints [30].
Perhaps the most exciting implication of our framework is that it suggests a paradigm shift: instead of talking about simple and complex ecosystems (as quantified, e.g., by community diversity), we should be distinguishing simple and complex properties of ecosystems. In the language we presented, this distinction is encoded in the shape of the Pareto front, as illustrated in Fig. 5. For some properties of interest, one might find that useful predictions require substantial microscopic detail (non-coarse-grainable; Fig. 5, red). However, all the properties in Fig. 3 exhibit the opposite behavior, whereby a comparatively coarse description may provide a “good-enough” predictive power at a fraction of the entropy of the fully microscopic description (coarse-grainable; Fig. 5, green). Our key result can be summarized as saying that even as community complexity increases, the prediction complexity of a given property need not do the same, and can even become simpler, as we quantitatively demonstrate.
An empirical confirmation of this behavior is very exciting because learning to harness emergent predictability could enable new methodologies for predicting and controlling complex ecosystems: methodologies for which community richness is an asset rather than a nuisance. Understanding the mechanistic origin of this effect, and developing a theoretical understanding for the conditions of its emergence, are important directions for future work.
@@
Our approach has limitations. Defining emergent simplicity requires specifying a class of models, and here, due to the limited size of the two available datasets, our choice (linear regression) was particularly simple. However, our theoretical framework is more general; with sufficient data, the natural generalization of the Pareto fronts in Figs. 1-3 would be to evaluate the predictive power of machine-learning models as a function of the dimension of the compressed (latent-space) representation of community composition. Understanding the extent to which emergent predictability depends on the choice of model class remains a question for future work. We should also note that in general, community richness (number of species) is a poor metric of community diversity, as it is overly sensitive to detection thresholds, and fails to account for the relatedness of community members [36]. This was not an issue for our work, as the synthetic communities considered here were assembled from a small pool of well-delimited taxa. However, for natural communities, this simple metric of diversity would need to be replaced by one of the better-suited metrics developed in the ecological literature [37, 38, 39].
@@
All simulations performed in MATLAB (Mathworks, Inc.). The associated code, data and scripts to reproduce all figures in this work are available at https://github.com/mtikhonov/emergentPredictability.
@@ Acknowledgements.
We thank S. Allesina, A. Goyal. K. Gowda, C. Holmes, S. Kaplan, S. Kuehn, R. McGee, A. Murugan, O. Venturelli and K. Wood for helpful discussions. This work was supported in part by the National Science Foundation grant PHY-2310746, and the Michigan Pioneer Fellows Program award to JM.
Supplementary information
.1 Quantifying description complexity, $I(\Psi)$
@@
It is important to emphasize that either choice preserves the phenomenon described in this work (compare main text Fig. 3 with Fig. S1; in either case, the Pareto fronts exhibit the shift indicative of emergent predictability). To explain our choice, we note that the entropy of $\Psi$ is the mutual information between the label of the coarsened class and the identity of the microscopic strain. Consider a scenario where two choices of two-group coarsenings are equally predictive, but one splits strains 50-50, while the other coarsening separates one strain from the rest of the taxa put together. If the latter is equally predictive, our definition identifies it as simpler, and thus preferable (same predictive power, lower complexity).
@@
Figure S1: Replotting result shown in Fig. 3 with description complexity axis $I(\Psi)$ measured in terms of number of classes of a given coarsened description $\Psi$.
.2 Simulations of standard ecological models
We investigate the concepts of emergent simplicity in two standard ecological models: a Generalized Lotka-Volterra (GLV) model and a Consumer Resource (CR) model. The GLV model we consider takes the following form:
$\dot{n}_{i}=\frac{r_{i}}{K_{i}}n_{i}\left(K_{i}-n_{i}-\sum_{j,j\neq i}a_{ij}n_{j}\right),$
which describes the abundance dynamics of each strain $i$ in terms of its growth rate $r_{i}$, carrying capacity $K_{i}$ and interactions with other strains encoded in $a_{ij}$. For simplicity, we set the growth rates for all strains to be the same $r_{i}=1$ since the key parameters determining equilibrium properties are the $K_{i}$ and interactions $a_{ij}$ [40]. We then adopt the approach common in large-$N$ theoretical ecology, drawing strain interactions $a_{ij}$ and carrying capacities $K_{i}$ randomly from specified distributions. Following refs. [41, 40], we use a normal distribution for both $K_{i}$ and $a_{ij}$. For the $K_{i}$, we set the mean to 1 and $\operatorname{std}(K_{i})=0.3$. For the interactions $a_{ij}$, the mean and spread of the distribution scale with the number of strains as $\mu/N$ and $\sigma/\sqrt{N}$, respectively. We set $\mu=1$ and $\sigma=0.8$ (modified from the “simplest” choice $\mu=1$ and $\sigma=1$ to facilitate the construction of a coexisting species pool). To obtain a strain library from which to subsample as illustrated in Fig. 1A, we first simulate the dynamics of a larger pool to find a set of coexisting strains (abundance $n_{i}>10^{-6}$). With our parameter choices, we found that a random set of 45 strains resulted in a community of 26 coexisting strains, similar to the high richness communities of the empirical datasets we consider in this work.
The CR model we choose to study describes the abundance dynamics of strains competing for a set of resources as follows:
$\dot{n}_{i}$ $=n_{i}\left(\sum_{\alpha}\sigma_{i\alpha}h_{\alpha}-\chi_{i}\right)$
$h_{\alpha}$ $=h_{\alpha}(T_{\alpha})\equiv\frac{b_{\alpha}}{1+T_{\alpha}/K_{\alpha}}$
$T_{\alpha}$ $\equiv\sum_{i}n_{i}\sigma_{i\alpha}.$
These are the same dynamics postulated in equation 1 of ref. [30] with slightly adjusted notation. Briefly, each strain $i$ is described as a binary phenotype $\sigma_{\alpha i}\in\{0,1\}$ indicating which resources $\alpha$ each strain can or cannot consume. The growth benefit of each resource $h_{\alpha}$ is depleted as the total demand $T_{\alpha}$ for resources increases with increasing abundance of competing strains. The strength of this depletion is set by parameters $b_{\alpha}$ and $K_{\alpha}$. For simplicity, we take $b_{\alpha}=1$ for all resources $\alpha$. Finally, each strain is also described by a growth cost, which we choose to model in the same way as ref. [42], $\chi_{i}=c+\sum_{\alpha}\lambda\sigma_{i\alpha}+\epsilon_{i}$. As in ref. [30], we set the baseline cost $c=0.1$ and the cost per resource consumed $\lambda=0.5$. Using the same approach as in studying the GLV model, we draw parameters $K_{\alpha}$, $\epsilon_{i}$ and $\sigma_{i\alpha}$ randomly. The “carrying capacities” $K_{\alpha}$ are Gaussian distributed with mean $10^{10}$ and standard deviation $10^{9}$. (Note that the scale of abundance is of course arbitrary, and could have been set to 1. The value of $10^{10}$, meant to evoke the large typical size of microbial populations, was chosen to visually underscore the difference between the main text Fig. 2C (plotting the absolute variance of an invader abundance) and the other panels, such as Fig. 2A, B showing the fraction of variance.) Random contributions to each strain’s cost $\epsilon_{i}$ are also Gaussian distributed with mean 0 and standard deviation of 0.01. Each strain phenotype $\vec{\sigma}_{i}$ is a random $L$-dimensional binary vector where each element represents the ability to consume a resource, and has probability $p=0.5$ to be 1. Just as we did in the GLV model above, we assemble a community from a larger pool of $M$ phenotypes, choosing $M=50$ and $L=40$ such that the number of coexisting strains is similar to the empirical datasets ($\simeq$25).
In both the GLV and CR models, we use the same protocol for our computational experiments, which follows the scheme outlined in Fig. 1A and mimics the laboratory experiments we consider in this work. Using the same distributions used to form the complete community, we randomly draw a focal strain to invade subsets of this community. Specifically, we assemble 1000 communities for each richness tier: low richness with 1 to 5 strains, mid richness with 11 to 15 strains, and high richness with 21 to 25 strains. After a specified amount of simulation time, we note the abundances of all community members, recording them as the “pre-invasion abundances”; these will be used as predictors. We then computationally “invade” the communities with the focal strain, and measure the post-invasion abundance of the focal strain after the same amount of simulation time has elapsed; this abundance is recorded as the response variable we seek to predict. Fig. S2 shows the resulting Pareto fronts inferred by our algorithm (see below for details) for 10 random invading strains. These are the Pareto fronts we average to make Fig. 2 in the main text.
Figure S2: Plotting the Pareto fronts for predicting post-invasion abundance of randomly drawn strains in a Generalized Lotka-Volterra (GLV) model (top row) and Consumer-Resource (CR) model (bottom row). The panels shown in Fig. 2 are averages over the 10 strains in each ecological model. See text for simulation protocol details.
.3 Description of empirical datasets
Scanning the literature for empirical studies that assemble communities from a fixed library of strains, we identified 9 candidate datasets for further consideration. These are listed in Table S1 below, cataloguing which datasets satisfy the criteria required for analysis within our framework.
Dataset Ecological Context $\geq 2$ Richness Tiers (yes/no) Number of Samples / Number of Strains (by Tier) Satisfies Inclusion Criteria?
[33] microbial yes 870/25 (low), 477/25 (mid), 180/25 (high) yes
[34] microbial yes 19433/14 (low), 346/14 (mid) yes
[43] microbial yes 21/17 (low), 30/17 (mid) not enough data
[44] grasslands yes 350/31 (low), 200/31 (mid), 50/31 (high) not enough data; not all strains were sampled in low and mid richness samples
[45, 46] grasslands yes 345/32 (low), 22/32 (mid) not enough mid richness samples
[47, 48] grasslands yes 200/60 (low), 40/60 (mid), 10/60 (high) not enough data
[49, 50, 51] grasslands no N/A too few richness levels sampled and not enough data
[52] grasslands no N/A plots either had all or one species present
[53] grasslands no N/A low richness only
Table S1: Catalogue of candidate published datasets that sample from a fixed library of strains. The last three columns indicate which datasets satisfy the required inclusion criteria for our framework to be applicable.
In summary, we determined that the datasets reported in [33] and [34] satisfy all requirements for applying our framework; namely, (1) enough community richness levels are sampled to clearly define at least two distinct richness tiers for comparison, and (2) enough communities are sampled within each richness tier to train and test linear regression models.
Clark et al. [33] investigated the production of fermentation end-products (butyrate, acetate, lactate, succinate) in synthetic gut communities consisting of 1 to 25 strain isolates. Final strain abundances were determined by first measuring sequenced read counts to compute relative abundances, and then calculating absolute abundances by multiplying by total $\mathrm{OD}_{600}$ of each community. These are the data use as inputs to our regression analysis for predicting community function (see below).
The experimental platform of Kehe et al. [34] was a microfluidic device containing various geometrical arrangments of microwells for high-throughput, parallel screening of community composition to function mapping. Each microwell array enables the grouping and merging of nanoliter droplets that contain different strain isolates from soil. Unlike Clark et al. [33], abundances of input strains were not measured; only the abundance of the focal strain H. Frisingense was tracked in arbitrary fluorescence units. Because of this, instead of using abundance as inputs to our framework presented in Fig. 1B, we use the presence ($n_{i}=1$) or absence ($n_{i}=0$) of strain droplets reported by Kehe etal as our inputs for predicting the final abundance of the focal strain. When we apply a given coarsening $\Psi$, we sum up the presence of strains that are grouped together for each sample community.
.4 Normalization of predictor and response variables
Experimentally, if we think about e.g. soil communities, it would be reasonable to define properties of interest $Y$ as measured in “per gram biomass.” In this case the overall biomass is normalized away, and this is the scenario we had in mind when calling $\Psi_{z}$ the “zero-information” description in Fig. 1C. This is equivalent to trying to predict the property of interest from relative abundances of taxa.
In lab experiments, where an inoculum is provided a fixed amount of resources, normalizing to biomass may not be necessary, since properties of interest can be defined in the absolute sense. (In fact, the resulting biomass could itself be a property of interest.) In this case, $\Psi_{z}$ is not actually “zero-information,” since knowing the overall biomass can be predictive (in which case the fraction of unexplained variance will be smaller than 1).
In our analysis, rather than using the actual abundances of species, we standardized all input variables $n$ to zero mean and unit variance. All coarsening schemes were then applied to these standardized variables. It is important to note that summing standardized abundances is not equivalent to standardizing summed abundances. However, the experimentally available numbers (16S reads) are only proxies for the absolute abundances anyway. The proportionality factor includes the number of 16S copies on the genome, which differs between species, and an PCR amplification bias, which is unknown and often substantial [54, 55]. Furthermore, even if absolute abundances were known, it is not clear if the summation rule implementing the intuition of $\tilde{n}$ as putative “functional classes” should be in units of biomass, in units of metabolic activity, or weighted in some other way.
@@
Thus, although this intuition of $\tilde{n}$ as the combined abundance of a putative functional class makes it easiest to explain our approach, any practical implementation on the available experimental data would implicitly include an unavoidable species-specific factor in front of every abundance value. Given this, we chose to fix these factors by simply standardizing input variables to unit variance, which is generally good practice with regression models, and came with the added benefit of improving performance. That said, we did also confirm that using raw OD-scaled 16S counts without standardization retains our conclusion (Fig. S3).
@@
The standardization step would certainly complicate attempts at mechanistic interpretation of the low-complexity descriptions discovered by our framework (lying along the Pareto front). In general, this strategy—discover a coarse-graining that is predictive, and then gain mechanistic insight by understanding why it works—is one of the promising applications of a framework like ours, or feature selection frameworks more generally. However, in this work, we focus on demonstrating the existence of diversity-aided emergent simplicity. Interpreting the variables is left for future work.
.5 Training and testing of linear regression models

### v3

@@ III Discussion
The promise of emergent simplicity is to enable predictions that work because of diversity, not despite it. However, so far, the field was unable to reach a consensus on whether emergent simplicity is real, whether it is surprising, and whether it is useful. Answering the first question requires a quantitative definition; the second requires a null model for comparison. For the third, we need an explicit focus on prediction rather than reproducibility, and a clear understanding of whether emergent simplicity is a biological statement, or a statistical one. Indeed, predictions that are a direct mathematical consequence of the number of species being large, independent of the underlying biology, are unlikely to be sufficiently nontrivial to be useful in practice.
@@
Here, we developed an information-theoretic framework to quantify the notion of emergent simplicity, and applied this framework to data from two separate experiments on microbial communities assembled in the lab. Our approach enabled us to define and distinguish coarse-grainability and emergent predictability, and to contrast both of these phenomena with “emergent reproducibility” (self-averaging) considered in much previous work.
All empirical properties we considered showed a clear signature of coarse-grainability. However, as we discussed, this is not necessarily surprising since for any real ecosystem, at least some degree of coarse-grainability is expected. This does not mean that coarse-grainability is uninteresting. Some empirical reports indicate a truly striking extent of coarse-grainability [31], and it is not clear that this can be fully explained by the relatedness of community members, especially given the extensive, functionally relevant diversity that is known to persist even at the level of strains [23]. A clear quantitative metric of coarse-grainability, as presented here, should help investigate this in future work.
Our main findings concerned emergent predictability, whereby coarser compositional descriptions become more predictive at higher richness. Our key result is that this phenomenon is robustly supported by empirical data; yet is absent from standard models of theoretical ecology, making it highly non-trivial and surprising.
In four of the five cases (measurements of Clark et al.), the phenomenon appeared to be linked to the dominant structuring role of pH in these communities. This observation echoes other recent examples where coarse-grained predictions are enabled by general physiological [18], metabolic [71], or toxicity-mediated constraints [13]. We described a mechanism explaining how structuring factors of this kind may acquire a higher explanatory power at higher diversity. We illustrated it in a minimal model demonstrating how emergent predictability can arise from a physiology-mediated feedback with a collective environmental variable.
Perhaps the most exciting implication of our analysis is that it reconciles the generality of the central limit theorem with the possibility of nontrivial predictions. Indeed, under the mechanism we proposed, the diversity-dependent simplification is ultimately still a form of self-averaging – but one that leads to a stronger collapse onto a lower-dimensional subspace governed by metabolic or physiological constraints (Fig. 5F). Understanding this behavior will require extending the theoretical investigations of large-$N$ ecosystems beyond the Gaussian-interaction regimes [15, 40, 4], and integrating these ongoing efforts with new models capable of capturing the role of such structuring factors.
@@
Our approach has limitations. Our definitions of emergent simplicity required specifying a class of models, and here, due to the limited size of the available datasets, our choice (linear regression) was particularly simple. However, our theoretical framework is more general; with sufficient data, the natural generalization of the Pareto fronts in Figs. 1, 3-5 would be to evaluate the predictive power of machine learning models as a function of the dimension of the compressed (latent-space) representation of community composition, similar to the approach of Ref. [54]. Understanding the extent to which emergent predictability depends on the choice of model class remains a question for future work. We should also note that in general, community richness (number of species) is a poor metric of community diversity, as it is overly sensitive to detection thresholds, and fails to account for the relatedness of community members [53]. This was not an issue for our work, as the synthetic communities considered here were assembled from a small pool of well-delimited taxa. However, for natural communities, this simple metric of diversity would need to be replaced by one of the better-suited metrics developed in the ecological literature [29, 44, 14].
@@
All simulations performed in MATLAB (Mathworks, Inc.). The associated code, data and scripts to reproduce all figures in this work are publicly available [46].
@@ Acknowledgements.
We thank S. Allesina, A. Goyal. K. Gowda, T. GrandPre, C. Holmes, S. Kaplan, S. Kuehn, R. McGee, A. Murugan, O. Venturelli and K. Wood for helpful discussions. MT acknowledges a CAREER award from the National Science Foundation (PHY-2340791). This work was also supported by the National Science Foundation grant PHY-2310746, and the Michigan Pioneer Fellows Program award to JM.
Supplementary Text
Appendix A Data selection: Other empirical datasets available
Scanning the literature for empirical studies that assemble communities from a fixed library of strains, we identified 9 candidates for further consideration. These are listed in Table S1, and identify the reasons why only the datasets of Refs. [11] and [33] satisfy all requirements for applying our framework; namely, (1) enough community richness levels are sampled to clearly define at least two distinct richness tiers for comparison, and (2) enough communities are sampled within each richness tier to train and test linear regression models.
Appendix B Constructing the Pareto fronts – Additional details
B.1 Quantifying description complexity, $I(\Psi)$
@@
It is important to emphasize that either choice preserves the phenomenon described in this work (compare main text Fig. 4 with Fig. S1; in either case, the Pareto fronts exhibit the shift indicative of emergent predictability). To explain our choice, we note that the entropy of $\Psi$ is the mutual information between the label of the coarsened class and the identity of the microscopic strain. Consider a scenario where two choices of two-group coarsenings are equally predictive, but one splits strains 50-50, while the other coarsening separates one strain from the rest of the taxa put together. If the latter is equally predictive, our definition identifies it as being simpler, and thus preferable (same predictive power, lower complexity).
@@
B.2 Normalization choices
B.2.1 Normalizing the response variable (the community-level property $Y$)
The question of quantifying the complexity of description, discussed above, is closely tied to the choice of the property of interest, and in particular, its normalization. The easiest way to see this is to observe that with our definition of $I(\Psi)$, the one-group description $\Psi_{B}$ (all taxa grouped together into a single category) has complexity $I(\Psi_{B})=0$. Technically, this is not a zero-information description: if we think in terms of raw abundances of taxa, summing them is essentially the total community biomass, which can be informative. However, in many situations (say, if studying process rates in natural soil), properties of interest scale with biomass, and the quantity most natural to predict is the value normalized per unit biomass. In such cases, our choice of $I(\Psi)$ is particularly appropriate. In contrast, in lab experiments where an inoculum is provided a fixed amount of resources, normalizing to biomass may not be necessary or desired, since properties of interest are meaningfully defined in the absolute sense — in fact, the final biomass of the assembled community could itself be a property of interest.
Fortuitously, in the empirical datasets used here, the one-group coarsening (to which we colloquially refer as the “biomass-only description”) explained almost no variance (never exceeding $15\%$ across all richness tiers and measured properties). This simplified our presentation, allowing us to declare that we were primarily focused on scenarios where the one-group coarsening explained no variance, such that we could equate $I(\Psi_{B})=0$ with “having no compositional information.” In the interest of conciseness, in the main text it was easiest to frame this as a decision to limit consideration to only a subset of community properties – those for which this statement is true. However, now is a good time to clarify that properties for which biomass alone explains a significant fraction of the variance are also compatible with our approach / framework. The generalization is straightforward: instead of the fraction of total variance explained, one should consider the residual variance after regressing on the biomass. With this approach, the Pareto fronts will start at $\mathrm{FVU}=1$ by definition, which would in fact simplify some downstream analyses. However, defining the $y$ axis of our diagrams in this way would make them appear rather less intuitive on first encounter. To summarize, limiting consideration to properties that are not explainable by biomass is a choice made for ease of presentation, not a limitation of our framework.
One minor consequence of this choice was for the computational invasion experiments (GLV & CRM), where making this analysis comparable with our empirical curves required us to exclude the instances ($\approx 15\%$ of cases) where the combined standardized abundances could explain >>10% of total variance in simulated “community invadability” data at high richness. Had this variance been defined as the residual variance after regressing on the biomass-only prediction, applying such exclusion criteria would not have been necessary. However, the advantage offered by greater simplicity of presentation of our key concepts outweighed this consideration.
B.2.2 Normalizing the predictor variables
In our analysis, the input variables supplied as “abundance matrix” $n$ were first standardized to zero mean and unit variance. All coarsening schemes were then applied to these standardized variables.
To understand this choice, it is important to note some limitations of the datasets. The study of Kehe et al. [33] did not directly measure the full composition of the final assembled communities: only the abundance of the focal strain H. Frisingense was tracked (in arbitrary fluorescence units). The available compositional information is limited to the makeup (binary presence/absence) of the set of droplets combined by the authors’ microfluidic platform to form the community initially (each droplet carrying individuals of a given strain). These binary variables are proxies for the initial abundances of community members, and in this sense, they still satisfy the criterion of Fig. 1 of “following a consistent assembly protocol from a known compositional state.” However, the relationship between this binary $n$ matrix and the actual abundance values is only approximate, and includes unknown proportionality factors.
In the study of Clark et al. [11], the composition of final assembled communities was profiled more directly, using 16S sequencing (recording relative abundances) as well as recording the $\mathrm{OD}_{600}$ for each community. However, even here, the best estimate of the absolute abundances (multiplying the relative fractions of 16S counts by the recorded community OD) is still only a proxy. The proportionality factor includes the number of 16S copies on the genome, which differs between species, and a PCR amplification bias, which is unknown and often substantial [35, 43]. Because of this, we adopted an approach where the input abundances were standardized to zero mean and unit variance (across samples) before training any regression models. For consistency, the same standardization procedure was also applied to Kehe et al. data.
@@ B.2.2 Normalizing the predictor variables
It is important to note that summing standardized abundances is not equivalent to standardizing summed abundances. However, even if absolute abundances were known, it is not clear if the biologically meaningful summation should be best implemented in units of biomass, in units of metabolic activity, or weighted in some other way. Thus, although this intuition of $\tilde{n}$ as the combined abundance of a putative functional class makes it easiest to explain our approach, any practical implementation on the available experimental data would implicitly include an unavoidable species-specific factor in front of every abundance value. Given this, we chose to fix these factors by simply standardizing input variables to unit variance, which is generally good practice with regression models, and came with the added benefit of improving performance. That said, we did also confirm that omitting the standardization step retains our conclusion (Fig. S2).
@@
This standardization of abundances would certainly complicate attempts at mechanistic interpretation of the low-complexity descriptions discovered by our framework (lying along the Pareto front). In general, this strategy—discover a coarse-graining that is predictive, and then gain mechanistic insight by understanding why it works—is one of the promising applications of a framework like ours, or feature selection frameworks more generally. However, in this work, we focus on demonstrating the existence of diversity-aided emergent predictability. Interpreting the high-performance groupings is left for future work.
B.3 Convergence of the Pareto search algorithm
We found the convergence of our Metropolis algorithm to be most efficient if the search is performed leveraging all of the data, and evaluates a candidate grouping using its in-sample explanatory power. Indeed, if using out-of-sample performance, one must use a very large number of cross-validation folds. Otherwise, the stochasticity of cross-validation (the variability induced by drawing different subsamplings of data) can mask true differences in performance, especially when applied to relatively small datasets like here, so that the decision whether to accept a given Metropolis move can become effectively random, impeding convergence. This large number of cross-validation folds sampled at each evaluation step would increase the runtime by orders of magnitude. Thus, in our approach, during the search phase only, performance of a candidate grouping is assessed using the in-sample FVU. After convergence, the performance of each grouping identified as belonging to the Pareto front is reassessed using out-of-sample cross-validation. This methodology was previously shown to ensure good convergence on synthetic datasets with known ground truth [74], and performed well in our case as well (see section B.5, “Robustness tests”).
B.4 Training and testing of linear regression models